Could someone please help me with this problem. I need to find the equation of the line parallel to 2y = -x + 3 which passes through (0,4)

My solution was to simplify 2y = -x + 3
So I tried 2y/2 = -x/2 + 3/2
so I came up with y = -x/2 + 3/2 but I am yucky with fractions. Could someone tell me if I am on the right page or not?

Answer by stanbon(75887)

find the equation of the line parallel to 2y = -x + 3 which passes through (0,4)
-------------------------
The given equation has a slope:
y = (-1/2)x + (3/2)
slope = -1/2
-------------------------------
Any line parallel to it must have slope = -1/2
------------------------------
Form of the line is y = mx+b
You have y,x, and m; solve for b.
4 = (-1/2)*0 + b
b = 4
--------------
EQUATION:
y = (-1/2)x + 4
================================
Cheers,
Stan H.